An object with a mass of 3.50 kg moves in a straight line on a frictionless horizontal surface with a speed of 2.0 m/s. Find the kinetic energy and the net work done on the object to make it reach this speed if it started from rest.


1
Expert's answer
2022-02-10T08:37:03-0500

Given:

"v_f=2.0\rm\: m/s"

"v_i=0\rm\: m/s"

"m=3.50\rm\: kg"


The final kinetic energy

"E_{kf}=\frac{mv_f^2}{2}=\frac{3.50*2.0^2}{2}=7.0\:\rm J"

The work done

"W=E_{kf}-E_{ki}=7.0-0=7.0\:\rm J"